40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5147     Entry time: Tue Aug 9 02:03:16 2011
Author: kiwamu 
Type: Summary 
Category: SUS 
Subject: summary of today's work on ETMY 

[Rana / Jenne / Kiwamu]

The ETMY suspension tower is currently sitting on the north side of the table for some inspections.

The adjustment of the OSEMs is ongoing.

 

(What we did)

  + Taken out two oplev mirrors, Jamie's windmill and a lemo patch panel.

  + Put some pieces of metal as makers for the original place

  + Put some makers on the distance of  dLY = -25.49 cm = -10.04 inch from the original place (see the 40m wiki).

     The minus sign means it will move away from the vertex.

  + Brought the ETMY suspension tower to the north side to do some inspections

  + Did some inspections by taking the noise spectra (#5141)

  + Adjusted the OSEM range and brought the magnets on the center of the OSEM holders by rotating and translating the OSEMs

  + During the work we found the proper PIT and YAW gains were about -5, which are the opposite sign from what they used to be.

  + Trying to minimize the cross couplings

JD: There is still some funny business going on, like perhaps the LR magnet isn't quite in the OSEM beam.  We leave the optic free swinging, and will continue to investigate in the morning.

ELOG V3.1.3-